# Neuronal Spike Trains as Functional-Analytic Distributions: Representation, Analysis, and Significance

### v1

@@ Abstract
The action potential constitutes the digital component of the signaling dynamics of neurons. But the biophysical nature of the full time course of the action potential associated with changes in membrane potential is fundamentally and mathematically distinct from its representation as a discrete set of events that encode when action potentials triggered in a collection spike trains. In this paper, we rigorously explore from first principles the transition and modeling from the standard biophysical picture of a single action potential to its representation as a spike in a spike train. In particular, we adopt a functional-analytic framework, using Schwartz distribution theory to represent spike trains as generalized Dirac delta functions acting on smooth test functions. We then show how and why this representation transcends a purely descriptive formalism to support deep downstream analysis and modeling of spike train neural dynamics in a mathematically consistent way.
@@ 1 Introduction
An action potential, as a dynamic transient change in the membrane potential of a neuron, has the shape and kinetics it has because of the underlying biophysics of the change in membrane permeability (conductance) to (primarily) $Na^{+}$ and $K^{+}$ ion channels. The conditions responsible for the diffusion (concentration) and electrochemical driving forces that move ions across the membrane are necessarily set up before an action potential is triggered, resulting in the instantaneous passive flux of $Na^{+}$ and $K^{+}$ ions acting as charge carriers responsible for the currents and resultant transient changes in membrane potential we call an action potential.
We assume the reader is familiar with (at least the qualitative) thermodynamic and biophysical description of the action potential summarized in the preceding paragraph. To varying degrees of detail, any standard textbook in neurophysiology can be consulted. Vander’s Human Physiology is particularly good at an introductory undergraduate level [widmaier2023vanders], while Kendel et al. is the standard graduate-level textbook in the field [kandel2021principles]. Although there are many other superb resources, both online and in print.
Under normal physiological conditions, the shape of the standard neuronal action potential is stereotyped and unchanging. Every time an action potential is triggered, it proceeds to completion uninterupted in exactly the same way with the same time course, independent of anything else the neuron is doing. As alluded to above, this is because the biophysical conditions responsible for the processes that mitigate the changes in membrane potential are set a priori. From an information-theoretic perspective, the action potential constitutes the digital component of neurons’ signaling dynamics. The information it contributes to the system is in the timing of its occurrence, not in its shape, amplitude, or time course.
@@
Figure 1: The often shown but inacurate picture of an action potential, which is a stereotyped change in the membrane potential due to underlying biophysics, compressed to a spike intended to represent it.
@@
This immediately creates a disconnect with the mathematical machinery we typically use to describe and model physical systems. Differential equations, filters, convolutions, and dynamical systems are built to operate on functions of time: objects that assign a value to every instant. A list of isolated time points does not fit naturally into this framework. Critically, we need a way to mathematically model spike trains that maintains functional consistency. We need a description of spike trains that represents them for what they are, but in a way that supports interaction with functions and operations that model physical processes that are very different mathematically from the description of a spike train.
In this paper, we rigorously explore from first principles the transition and modeling from the standard biophysical picture of a single action potential to its representation as a spike in a spike train. In particular, we adopt a functional-analytic framework, using Schwartz distribution theory to represent spike trains as generalized Dirac delta functions acting on smooth test functions. We then show how and why this representation transcends a purely descriptive formalism to support deep downstream analysis and modeling of spike train neural dynamics in a mathematically consistent way.
@@
The ideas and content in this paper draw on several decades of work in theoretical and computational neuroscience and neural dynamics. Representative treatments ranging from early point-process formulations of spike trains to modern dynamical systems and network models can be found in [DayanAbbott2001], [GerstnerKistler2002], [GerstnerKistlerNaudPaninski2014], [Tuckwell1988], [PerkelGersteinMoore1967], [Knight1972], [BrunelHakim1999], and [ErmentroutTerman2010], among many others. See also the references within these cited works.
@@
The primary contribution of the present work is not the introduction of new neurophysiological mechanisms or descriptions, but a first-principles reconstruction of these concepts within a unified framework grounded in functional analysis and distribution theory [botelho2025, brayman2024, aksoy2024, muscat2024, clason2020]. By making the underlying mathematical structure explicit, the paper clarifies the sense in which spike trains are treated as distributions: mathematical objects defined not as functions or by their height at a single point, but by their interactions as unique events with smooth test functions that act as ’probes’. This framework provides a rigorous basis for treating spikes as mathematical objects that isolate timing and weight as the primary variables driving neural dynamics while maintaining consistency with neurophysiologically meaningful dynamical systems.
The intent is to provide a self-contained, technically precise reference that bridges what is often a communication barrier between the physiological and mathematical perspectives. It offers readers a rigorous foundation for engaging with and extending the extensive literature on spike-based neural dynamics.
@@ 2 The Representational Objective
To define a spike train, we need to define an object $s_{i}$ such that any physically meaningful operation we do with it can be expressed as an integral or sum paired with a test function $\varphi(t)$ that measures the effect of $s_{i}$ on the operation. A valid test function will need to satisfy a number of formal mathematical properties and will be of practical neurophysiological utility, as in the examples introduced in the previous paragraph.
@@
Equation 2.1 is then the operational requirement for a spike train. A set $s_{i}$ paired with a test function $\varphi(t)$ such that $s_{i}$ is defined by how it samples and evaluates the summed effect of $\varphi(t)$ only at specific times.
@@ 3 Why the Dirac delta is not a classical (Lebesgue) function
Almost immediately, one gets a sense that this is not a function in the ordinary sense. By an ‘ordinary’ or classical function, here, we mean an integrable Lebesgue function. A function that assigns a real number to (almost) every time $t$, in a way that is well-defined pointwise and compatible with Lebesgue integration. Formally, this means that $f:\mathbb{R}\to\mathbb{R}$ such that for each $t\in\mathbb{R}$ (or for almost every $t$, not including any discontinuities), the value $f(t)$ exists as a real number. This type of function is measurable with respect to the Lebesgue measure, so the integration of the function $\int f(t)\,dt$ makes sense and can be defined. (There are other conditions and implications, but for our purposes here, this is sufficient to make our point.)
@@ 4.1 Function vs functional
$F(\varphi):=\int_{a}^{b}\varphi(t)\,dt$ $\text{integrated on the closed intervarl }[a,b]$ (4.1b)
@@
The Dirac delta is a mathematical object (a functional) that takes a test function $\varphi(t)$ and returns a number, explicitly, the value of $\varphi(t)$ at the point $a$. Notice how the pairing $\langle\delta(t-a),\varphi\rangle$ produces a number, the value of the test function at $t=a$, and satisfies the operational requirement set up in Equation 2.1. This define the Dirac delta for the single-valued case, but not (yet) a spike train $s_{i}$, which we work up to below.
@@ 4.2 Linearitiy of functionals
Similarly, the Direct delta is linear because it satisfies
@@ 4.3 The space of test functions: Definition of support
Test functions are smooth, well-behaved functions that probe the distribution. We use test functions to ask questions and observe the effects of the distribution on them. Schwartz distributions, like the Dirac delta, are not defined pointwise. They do not evaluate an input that maps to an output like a regular function (see the discussion above in Section 3.) They only make sense in context with how they interact with test functions.
Test functions $\varphi(t)$ are not biophysically measurable or derived variables like the membrane potential, synaptic currents, or ionic conductances. They are any smooth time-localized filter that can be properly defined that can be used to interrogate the distribution. Explicitly, in our case, the distribution being constructed is a spike train encoded by Dirac deltas. We need to restrict the space of test functions to only those with these properties because, for arbitrary functions, evaluation at a point may not be defined, differentiation may be undefined, and convergence may not hold. All of which we need in the interaction between a functional and a test function. Below, we rigorously formalize these requirements.
To model a spike train, we can restrict the space of test functions to $C_{c}^{\infty}$, the space of smooth functions with compact support on $\mathbb{R}$. $C$ implies continuous functions, while $C^{\infty}$ implies continuous functions for which all derivatives exist. In other words, they are infinitely differentiable. This makes the space of test functions not just continuous but smooth, a stronger condition than continuity; i.e., all smooth functions are continuous, but not all continuous functions are smooth. Formally,
@@
The subscript in $C_{c}^{\infty}$ means the set of smooth functions that are compactly supported, meaning if $\varphi(t)\in C_{c}^{\infty}\rightarrow\varphi(t)=0$ outside some finite interval. Understanding what the support of a function is, and how it differs between ordinary (Lebesgue) functions and Schwartz distributions, is not just a technical mathematical detail; it is at the core of understanding why the Dirac delta works as a model of spike trains.
The requirement for test functions to be smooth ($C^{\infty}$) is a strategic choice that shifts the burden of differentiation. Because many distributions (like the spike train $s_{i}$) are singular events and lack classical derivatives, distribution theory transfers the operation of differentiation onto the test function. Because $\varphi(t)$is infinitely differentiable, the definition of differentiation of the spike train is dependent on the derivatives of the test function. This ensures that the mathematical machinery never breaks down when encountering singularity associated with the events that make up a distributional like the spike train. We discuss this in some detail in Section 7.4 at the end of the paper.
We will first define the support of ordinary functions, which applies to test functions $\varphi(t)$. We will then show why that definition fails for distributions, and finally, how we can adjust the definition of support to accommodate distributions. What will come out of this exercise is an understanding of what it means for Dirac deltas to represent and encode spikes in a spike train, and why the concept of Schwartz distributions only makes sense in context with their effects on test functions.
@@
There is a subtle yet critical notion here that can easily be missed. This definition relies on pointwise values. The support is defined as a subset of the function’s domain (the subset where it is nonzero), and includes boundary points wherever the function is nonzero arbitrarily close to them. This means that support is not defined by where a function is nonzero at that exact point. It is defined by where it is nonzero arbitrarily close to that point. Or put another way, a point in the domain of $f(t)$ belongs to the support if the function cannot be made identically zero in any arbitrarily small neighborhood around that point. This definition is designed to capture the notion of local influence, not pointwise values of $f(t)$ for some value of $t$. If this was not the case, and we attempted to define support on a pointwise basis for all nonzero values, i.e., $\{t:f(t)\neq 0\}$, then boudary points would be arbitrarily excluded, since any nonzero value of $f(t)$ could be made to approach the boundary point as close as we like. There would be no way to non-arbitrarily define where the function goes from being nonzero to zero. This would produce downstream inconsistencies. For example, integration by parts would fail. And, most importantly for us in this paper, the definition of support would not generalize to Schwartz distributions.
@@
If $f(t)\neq 0$ on the open interval $(0,1)$, and $f(0)=f(1)=0$, the support would still be the closed interval $[0,1]$, i.e., $\mathrm{supp}(f)=[a,b]$. The nonzero set in this case is $\{t:f(t)\neq 0\}=(0,1)$, but the closure of the open set is the closed set $[0,1]$ because every neighborhood of 0 contains points $t>0$ (to the right of 0) even though $f(0)=0$. And every neighborhood of $1$ contains points $t<1$ (to the left of 1) even though $f(1)=0$. So the points 0 and $1$ are limit points of the nonzero open set $(0,1)$, and are therefore included in the support of the function.
But if $f(t)\neq 0$ for $f(0)$ and $f(1)$, then it is nonzero only at isolated points, and differs from the the zero function $g(t)0$ $\forall t\in\mathbb{R}$ on a finite set (like we described above). Such a function has a Lebesgue support equal to the empty set. This is why attempting to use the Lebesgue definition of support fails for the Dirac delta and, more generally, for Schwartz distributions. Instead, the support of a distribution is defined operationally by how it acts on test functions.
@@ 4.4 The space of test functions: Continuity of functionals
$\varphi_{n}\to\varphi\quad\Longrightarrow\quad F(\varphi_{n})\to F(\varphi)$
@@
$\sup_{t}|D^{k}(\varphi_{n}-\varphi)(t)|\to 0\quad\text{for all }k$
and the supports remain bounded in a compact set, i.e., functions do not explode to infinity. If these conditions are met, it ensures that any $\varphi_{n}$ is indistinguishable from $\varphi$ even under arbitrarily fine local inspection. Note that ’sup’ here is not ’support’ but ’supremum’, meaning least upper bound, in this case taken over the entire variable $t$, which is why we write $\sup_{t}$. Notaionally, we are also using $D^{k}\equiv\frac{d^{k}\varphi}{dt^{k}}$. So writing $\sup_{t}|D^{k}(\varphi_{n}-\varphi)(t)|$ means the largest value ($\equiv$ the smallest possible upper bound) of all the derivatives of $(\varphi_{n}-\varphi)(t)$ as $t$ ranges over all real numbers, i.e., $t\in\mathbb{R}$.
@@
implying that the Dirac delta is, as required, a continous functional.

### v2

@@ Abstract
The action potential constitutes the digital component of the signaling dynamics of neurons. But the biophysical nature of the full-time course of the action potential associated with changes in membrane potential is mathematically distinct from its representation as a discrete set of events that encode when action potentials are triggered in a collection of spike trains. In this paper, we develop from first principles a unified functional-analytic framework for neuronal spike trains, grounded in Schwartz distribution theory. We show how this representation provides an exact operational calculus for convolution, distributional differentiation, and distributional support, which enables closed-form analysis of spike train dynamics without discretization, rate approximation, or smoothing. We then analyze the framework in the context of a two-neuron reciprocal circuit with propagation latencies and refractoriness, deriving exact results for synaptic drive, spike timing sensitivity, and causal admissibility of inputs, quantities that are either ill-defined or require approximation in conventional treatments.
@@ 1 Introduction
A neuronal action potential, as a dynamic transient change in the membrane potential of a neuron, has the shape and kinetics it has because of the underlying biophysics of the change in membrane permeability (conductance) to (primarily) $Na^{+}$ and $K^{+}$ ion channels. The conditions responsible for the diffusion (concentration) and electrochemical driving forces that move ions across the membrane are necessarily set up before an action potential is triggered, resulting in the instantaneous passive flux of $Na^{+}$ and $K^{+}$ ions acting as charge carriers responsible for the currents and resultant transient changes in membrane potential we call an action potential. For foundational references and background on the physiology and biophysics of the action potential, Vander’s Human Physiology is particularly good at an introductory undergraduate level [widmaier2023vanders], while Kendel et al. is the standard graduate-level textbook in the field [kandel2021principles]. Although there are many other superb resources, both online and in print.
Under normal physiological conditions, the shape of the neuronal action potential is stereotyped and unchanging. Every time an action potential is triggered, it proceeds to completion uninterrupted in exactly the same way with the same time course, independent of anything else the neuron is doing. As alluded to above, this is because the biophysical conditions responsible for the processes that mitigate the changes in membrane potential are set a priori. From an information-theoretic perspective, the action potential constitutes the digital component of a neuron’s signaling dynamics. The information it contributes to the system is in the timing of its occurrence, not in its shape, amplitude, or time course.
@@
Figure 1: The often shown but inacurate picture of an action potential, which is a stereotyped change in the membrane potential due to the underlying biophysics, compressed into a spike intended to represent it.
@@
This immediately creates a disconnect with the mathematical machinery we typically use to describe and model physical systems. Differential equations, filters, convolutions, and dynamical systems are built to operate on functions of time, objects that assign a value to every instant. A list of isolated time points does not fit naturally into this framework. Critically, we need a way to mathematically model spike trains that maintains functional consistency. We need a description of spike trains that represents them for what they are, but in a way that supports interaction with functions and operations that model physical processes that are very different mathematically from the description of a spike train.
In this paper, we develop from first principles a unified functional-analytic framework for neuronal spike trains, grounded in Schwartz distribution theory [botelho2025, brayman2024, aksoy2024, muscat2024, clason2020]. Spikes are represented as generalized functions defined by their action on smooth test functions rather than by pointwise values. We show that this representation is not merely a descriptive formalism, but rather, it provides an exact operational calculus for convolution, distributional differentiation, and distributional support, which enables closed-form analysis of spike train dynamics without discretization, rate approximation, or smoothing. We demonstrate the analytical power of this framework on a two-neuron reciprocal circuit with discrete signaling event propagation latencies and refractoriness, deriving exact results for synaptic drive, spike timing sensitivity, and causal admissibility of inputs, quantities that are either ill-defined or require approximation in conventional treatments.
@@
The ideas and concepts we bring together and build on draw on several decades of work in theoretical and computational neuroscience and neural dynamics. Representative treatments ranging from early point-process formulations of spike trains to modern dynamical systems and network models can be found in [dayan2001theoretical], [GerstnerKistler2002], [GerstnerKistlerNaudPaninski2014], [Tuckwell1988], [PerkelGersteinMoore1967], [Knight1972], [BrunelHakim1999], and [ErmentroutTerman2010], among many others. See also the references within these cited works.
@@
By making the underlying mathematical structure explicit, the paper clarifies the sense in which spike trains are treated as distributions: mathematical objects defined not as functions or by their height at a single point, but by their interactions as unique events with smooth test functions that act as ’probes’. This framework provides a rigorous basis for treating spikes as mathematical objects that isolate timing and weight as the primary variables driving neural dynamics, while maintaining consistency with neurophysiologically meaningful dynamical systems. The intent is to provide a self-contained, technically rigorous reference that bridges what is often a communication barrier between the physiological and mathematical perspectives. It offers readers a rigorous foundation for engaging with and extending the extensive literature on spike-based neural dynamics.
@@ 2 The Representational Objective
To define a spike train, we need to define an object $s_{i}$ such that any physically meaningful operation we do with it can be expressed as an integral or sum paired with a test function $\varphi(t)$ that measures the effect of $s_{i}$ on the operation. A valid test function will need to satisfy a number of formal mathematical properties and should be of practical neurophysiological utility, like the examples from the previous paragraph.
@@
Equation 2.1 is then the operational requirement for a spike train, where the brackets $\quad\langle\cdot,\cdot\rangle$ denote the action of the functional (spikes) on the test function. We elaborate on this paired duality notation in Section 4.1 below. It specifies a set $s_{i}$ paired with a test function $\varphi(t)$ such that $s_{i}$ is defined by how it samples and evaluates the summed effect of $\varphi(t)$ only at the specific times events (spikes) occur.
@@ 3 Why the Dirac delta is not a classical (Lebesgue) function
Almost immediately, one gets a sense that this is not a function in the ordinary sense. By an ‘ordinary’ or classical function, we mean an integrable Lebesgue function. A function that assigns a real number to (almost) every time $t$, in a way that is well-defined pointwise and compatible with Lebesgue integration. Formally, this means that $f:\mathbb{R}\to\mathbb{R}$ such that for each $t\in\mathbb{R}$ (or for almost every $t$, not including any discontinuities), the value $f(t)$ exists as a real number. This type of function is measurable with respect to the Lebesgue measure, so the integration of the function $\int f(t)\,dt$ makes sense and can be defined. (There are other conditions and implications, but for our purposes here, this is sufficient to make our point.)
@@ 4.1 Function vs functional
$F(\varphi):=\int_{a}^{b}\varphi(t)\,dt$ $\text{integrated on the closed interval }[a,b]$ (4.1b)
@@
The Dirac delta is a mathematical object (a functional) that takes a test function $\varphi(t)$ and returns a number, explicitly, the value of $\varphi(t)$ at the point $a$. Notice how the pairing $\langle\delta(t-a),\varphi\rangle$ produces a number, the value of the test function at $t=a$, and satisfies the operational requirement set up in Equation 2.1. This defines the Dirac delta for the single-valued case, but not (yet) a spike train $s_{i}$, which we work up to below.
In functional analysis, the standard notation $\langle T,\phi\rangle$ represents a duality pairing. It describes the interaction between two distinct mathematical objects: a functional $T$, in our case, the spike train, and a test function $\phi$. Although this notation resembles an inner product, such as the dot product between two vectors, in the functional-analytic context it reflects a fundamentally different operation. Instead of multiplying two functions point-by-point, this duality notation implies an operation, the explicit application of the rule defined by $T$ to the function $\phi$.
Formally, if $T\in\mathcal{D}^{\prime}(\mathbb{R})$ is a distribution and $\phi\in\mathcal{D}(\mathbb{R})$ is a test function, we write
$\langle T,\phi\rangle:=T(\phi)$
to denote the action of the continuous linear functional $T$ on $\phi$. $\mathcal{D}(\mathbb{R})$ is the space of infinitely differentiable test functions $\phi:\mathbb{R}\rightarrow\mathbb{R}$, while $\mathcal{D}^{\prime}(\mathbb{R})$ is the dual space of distributions on $\mathbb{R}$. In Sections 4.3 and 4.4 below we fully develop what these test functions are within the context of spike trains.
@@ 4.2 Linearity of functionals
Similarly, the Dirac delta is linear because it satisfies
@@ 4.3 The space of test functions: Definition of support
As introduced above in Section 3, test functions are smooth, well-behaved functions that probe a distribution. We observe the effects of the distribution on them. Schwartz distributions, like the Dirac delta, are not defined pointwise. They do not evaluate an input that maps to an output, and so only make sense in context with how they interact with test functions.
Test functions $\varphi(t)$ are not necessarily biophysically measurable or derived variables like the membrane potential, synaptic currents, or ionic conductances. They are any smooth time-localized filter capable of interrogating a distribution. Explicitly, in our case, the distribution being constructed is a spike train encoded by Dirac deltas. We need to restrict the space of test functions to only those with these properties because, for arbitrary functions, evaluation at a point may not be defined, differentiation may be undefined, and convergence may not hold. All of which we need in the interaction between a functional and a test function. Below, we rigorously formalize these requirements.
To model a spike train, we restrict the space of test functions to $C_{c}^{\infty}$, the space of smooth functions with compact support on $\mathbb{R}$. $C$ implies continuous functions, while $C^{\infty}$ implies continuous functions for which all derivatives exist, i.e., are infinitely differentiable. This makes the space of test functions not just continuous but smooth, a stronger condition than continuity; all smooth functions are continuous, but not all continuous functions are smooth. Formally,
@@
The subscript in $C_{c}^{\infty}$ means the set of smooth functions that are compactly supported, meaning if $\varphi(t)\in C_{c}^{\infty}\rightarrow\varphi(t)=0$ outside some finite interval.
The requirement for test functions to be smooth ($C^{\infty}$) is a strategic choice that shifts the burden of differentiation. Because many distributions (like the spike train $s_{i}$) are singular events and lack classical derivatives, distribution theory transfers the operation of differentiation onto the test function. Because $\varphi(t)$is infinitely differentiable, the definition of differentiation of the spike train is dependent on the derivatives of the test function. This ensures that the mathematical machinery never breaks down when encountering singularity associated with the events that make up a distribution like the spike train. We discuss this in some detail in Section 7.4 below.
@@
This definition relies on pointwise values. The support is defined as a subset of the function’s domain (the subset where it is nonzero), and includes boundary points wherever the function is nonzero arbitrarily close to them. This means that support is not defined by where a function is nonzero at that exact point. It is defined by where it is nonzero arbitrarily close to that point. A point in the domain of $f(t)$ belongs to the support if the function cannot be made identically zero in any arbitrarily small neighborhood around that point. This definition is designed to capture the notion of local influence, not pointwise values of $f(t)$ for some value of $t$. If this was not the case, and we attempted to define support on a pointwise basis for all nonzero values, i.e., $\{t:f(t)\neq 0\}$, then boundary points would be arbitrarily excluded, since any nonzero value of $f(t)$ could be made to approach the boundary point as close as we like.
@@ 4.4 The space of test functions: Continuity of functionals
$\varphi_{n}\to\varphi\quad\Longrightarrow\quad F(\varphi_{n})\to F(\varphi)$ (4.6a)
@@
$\sup_{t}|D^{k}(\varphi_{n}-\varphi)(t)|\to 0\quad\text{for all }k$ (4.6c)
and the supports remain bounded in a compact set, i.e., functions do not explode to infinity. If these conditions are met, it ensures that any $\varphi_{n}$ is indistinguishable from $\varphi$ even under arbitrarily local inspection. Note that ’sup’ here is not ’support’ but ’supremum’, meaning least upper bound, in this case taken over the entire variable $t$, which is why we write $\sup_{t}$. Notationally, we are also using $D^{k}\equiv\frac{d^{k}\varphi}{dt^{k}}$. So writing $\sup_{t}|D^{k}(\varphi_{n}-\varphi)(t)|$ means the largest value ($\equiv$ the smallest possible upper bound) of all the derivatives of $(\varphi_{n}-\varphi)(t)$ as $t$ ranges over all real numbers, i.e., $t\in\mathbb{R}$.
@@
implying that the Dirac delta is, as required, a continuous functional.
We end this section on an important technical note. The space of test functions $\mathcal{D}(\mathbb{R})$ is equipped with the standard canonical LF-topology, i.e., limit of Fréchet spaces. This topology is constructed as the strict inductive limit of the spaces $\mathcal{D}_{K}(\mathbb{R})$
[…]
where each subspace $\mathcal{D}_{K}(\mathbb{R})=\{\phi\in C^{\infty}(\mathbb{R}):\operatorname{supp}(\phi)\subset K\}$ consists of smooth functions supported within a compact set $K\subset\mathbb{R}$. The limit notation here is the inductive limit (or equivalently the direct limit). The space $\mathcal{D}(\mathbb{R})$ is the union of all the smaller spaces $\mathcal{D}_{K}(\mathbb{R})$, equipped with the strongest topology for which the inclusion maps $\mathcal{D}_{K}(\mathbb{R})\hookrightarrow\mathcal{D}(\mathbb{R})$ are continuous.
As a clarification point, notationally, we use $D^{k}$ in Equation 4.6 above for the $k$-th derivative operator, where $k$ is an integer. This should not be confused with the subspace $\mathcal{D}_{K}$ in Equation 4.7, where $K$ represents a compact set.
Mathematically, this construction is necessary because $\mathcal{D}(\mathbb{R})$ is too “large” to be metrizable. It cannot be described by a single distance metric or norm that simultaneously applies over the entire infinite real line. Instead, the LF-topology stitches together a family of simpler, metrizable spaces (Fréchet spaces) defined on finite time intervals.
Intuitively, this formalizes the physical constraint that experimental observations are always local. While a spike train may mathematically exist on an infinite timeline, any valid measurement that describes an action of the functional on a test function must occur within a finite observation window. Consequently, convergence in this space is defined locally, consisting of a sequence of test functions $\phi_{n}\to\phi$ in $\mathcal{D}(\mathbb{R})$ if and only if
  1. 1.
    There exists a common compact support (a fixed observation window) $K$ such that $\operatorname{supp}(\phi_{n})\subset K$ for all $n$, and
  2. 2.
    The functions and all their derivatives converge uniformly to $\phi$ within that window.
This formal (mathematical) characterization [brayman2024, clason2020] ensures that the operations of the spike train distribution remain stable and well-defined under the limits of biophysical recording.